On 1 July 2017 Miller Ltd acquired a 25% interest in Thomas Ltd for consideration of $73,000. At that date the equity of Thomas Ltd consisted of:

Share Capital                                140 000

Retained Earnings                          70 000

Asset Revaluation Surplus           12 000

                                                      222 000

All assets and liabilities of Thomas Ltd are recorded at fair value with the exception of inventory which was held at $5,000 below its fair value. The entire inventory was sold during the 2017-2018 financial year.

The following amounts represent the profit/(loss), dividends paid and asset revaluation surplus (ARS) balance by Thomas Ltd since acquisition:

Year ending 30/6/2018

$

Year ending 30/6/2019

$

Profit /(Loss) after tax

32 000

(20 000)

Dividend paid

12 000

7 000

ARS Balance

18 000

26 000

On 1 January 2018, Miller Ltd sold an item of equipment to Thomas Ltd for $18,000. The equipment had a cost of $30,000 and a carrying amount of $12,000 at the date of transfer. The equipment was estimated to have a further 4 years of useful life at the time of sale.

Tax rate is 30%.

Required:

Prepare the equity journals at 30 June 2018 and 30 June 2019 to account for the investment in Thomas Ltd in accordance with AASB 128, assuming Miller Ltd does not prepare consolidated financial statements. (10 marks). This question is based on equity accounting of corporate accounting course under Masters of accounting degree. If you do not know the answer, do not bother to ask for more info as the full question is given to you.
Answer #1

Caluclation of goodwill/capital reserve on the date of acquisition:

A) Cost of acquisition = 73000

B) Proportionate value of net assets

Share capital = 140000

Retained earnings = 70000

Asset revaluation surplus = 12000

Add: Inventory recorded at fair value = 5000

Total = 227000

Proportionate value of assets = 227000*25% = 56750

Goodwill on acquisition = 73000-56750 = 16250

It is given that the compnay is not preparing consolidated financial statements. Then the profit/ loss in the associate company will not be included in the books of accounts but the dividend received by the company will be adjusted against the goodwill amount which was created on the date of acquisition.

Date Particulars Debit Credit
30-06-2018 Bank A/c(12000*25%) 3000
To Dividend 3000
(Being dividend received from associate company)
30-06-2018 Dividend A/c 3000
To Goodwill A/c 3000
(Being dividend received from Thomson Ltd adjusted against Goodwill)
30-06-2019 Bank A/c (7000*25%) 1750
To dividend 1750
(Being dividend received from associate company)
30-06-2019 Dividend A/c 1750
To Goodwill 1750
(Being dividend received from Thomson Ltd adjusted against Goodwill)

If company adopts to consolidation, then the profit of associate need to be accounted.

Feel free to ask any doubts. Rate